Circular No. 6832

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998S IN NGC 3877 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report: “A spectrum taken by J. Huchra with the 1.5-m Tillinghast telescope on Mar. 5.3 UT confirms that SN 1998S is a narrowlined type-II supernova (type IIn; cf. IAUC 6830) but shows it to be quite unusual. There are Balmer emission lines with narrow peaks and broad wings superimposed on a very blue continuum. The full width at the base of Hα is 8200 km/s. What makes this type-IIn object stand out are the strengths of He II 468.6-nm and permitted N III/C III 464-nm, which are comparable to the flux of Hβ. These features are identical those seen in SN 1983K (Niemela et al. 1985, A.J. 289, 52), which was discovered two weeks before maximum light. There are also weaker He II lines seen at 541.1 and 569.5 nm. A broad emission feature peaking at 580.0 nm may be C IV, characteristic of Wolf-Rayet stars, and another broad (12-nm FWHM) emission at 710 nm is likely to be a blend of C II lines, also found in Wolf-Rayet stars. Narrow emission of He I at 706.5, 667.8, and 587.5 nm are detected, as well as weak [O III] at 436.3 nm. Na I absorption with an equivalent width of 0.07 nm is measured. The redshift of the narrow He I lines is 840 km/s. SN 1983K (in NGC 4699, redshift of 1318 km/s) reached a maximum of B = 12.4.”

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence

(OHP), write: “Optical CCD photometry of the counterpart of the transient Aql X-1, using a 15-min exposure obtained by A. Boselli (Laboratoire d’Astronomie Spatiale) with the 1.2-m OHP telescope on Mar. 6.2 UT, shows the object at V = 16.95 ± 0.02. This is already 0.5 mag brighter than the maximum level reached during the 1997 Aug.-Sept. outburst and shows that the current event, which started ∼ Feb. 27 (IAUC 6828), may reach levels comparable to those of the 1995 July-Aug. and 1997 Feb.-Mar. events. Monitoring will continue at OHP.”

SUPERNOVA 1997dq IN NGC 3810 Unfiltered magnitude estimate by T. Doi, Columbus, TX: Mar. 3.24

UT, 16.8 (accidental discovery).

1998 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green